# On the geometry of Kähler–Frobenius manifolds &\&& their classification

### v1

@@ 7.4. Hermitian–Einstein Space of Connections
We discuss the construction depicted throughout these sections. We prove that:
@@
Let $h$ be a Hermitian fibre–metric on $E$. Assume $\{s_{1},\cdots,s_{r}\}$ are linearly independent local holomorphic sections of $E$. Put $h_{i,\bar{j}}=h(s_{i},{\bar{s}_{j}})$.
With respect to the sections $\{s_{1},\cdots,s_{r}\}$ and the local coordinates $(z_{1},\cdots,z_{n})$ in $M$ the curvature of the Hermitian connection in the bundle $E$ is given by
@@
The quadruple $(E,M,h,ds^{2})$ is a Hermitian–Einstein vector bundle if the Ricci tensor is proportional to $h$:
@@
Let $\nabla_{0}$ be a Hermitian–Einstein connection. Let $\tilde{\mathscr{A}}$ be the affine space of connections. Any connection $\nabla\in\tilde{\mathscr{A}}$ can be written as $\nabla=\nabla_{0}+t\eta$, where $\eta$ is a section in $\Gamma(T^{*}M\otimes E^{*}\otimes E)$ and $t$ is a real parameter. The curvature $F_{\nabla}=d\nabla-\nabla\wedge\nabla$ of the Hermitian connection is a section of $EndE\otimes\Lambda^{2}T^{*}M$.
A Hermitian–Yang–Mills connection is a unitary connection $\nabla$ for the Hermitian metric $h$ which satisfies $F_{\nabla}^{0,2}=0$ i.e. vanishing of the curvature and the trace $tr$ of the curvature satisfies the following relation $tr(F_{\nabla})=\kappa Id_{E}.$
Given a Kähler metric on $M$, define the operation $tr:\Gamma(EndE\otimes T_{1,1}^{*}M)\to\Gamma(EndE)$, where $tr(F)=\sum g^{j\bar{k}}F_{aj\bar{k}}^{b}$.
We have a Hermitian–Yang–Mills if there exists a Hermitian metric $h$ for which the Hermitian curvature $F$ satisfies $tr(F)=\kappa Id$. If we take $E=TM$ to be the holomorphic tangent bundle and $h=g$ then the condition above becomes the familiar Kähler-Einstein condition (definition from Uhlenbeck–Yau).

### v2

@@ 7.4. Hermitian–Einstein Space of Connections
We discuss the construction depicted throughout these sections.
We prove that:
@@
Let $h$ be a Hermitian fibre–metric on $\pazocal{E}$. Assume $\{s_{1},\cdots,s_{r}\}$ are linearly independent local holomorphic sections of $\pazocal{E}$. Put $h_{i,\bar{j}}=h(s_{i},{\bar{s}_{j}})$.
With respect to the sections $\{s_{1},\cdots,s_{r}\}$ and the local coordinates $(z_{1},\cdots,z_{n})$ in $M$ the curvature of the Hermitian connection in the bundle $\pazocal{E}$ is given by
@@
The quadruple $(\mathscr{E},M,h,ds^{2})$ is a Hermitian–Einstein vector bundle if the Ricci tensor is proportional to $h$:
@@
Let $\nabla_{0}$ be a Hermitian–Einstein connection. Let $\tilde{\mathscr{A}}$ be the affine space of connections. Any connection $\nabla\in\tilde{\mathscr{A}}$ can be written as $\nabla=\nabla_{0}+t\eta$, where $\eta$ is a section in $\Gamma(T^{*}M\otimes\pazocal{E}^{*}\otimes\pazocal{E})$ and $t$ is a real parameter. The curvature $F_{\nabla}=d\nabla-\nabla\wedge\nabla$ of the Hermitian connection is a section of $End\pazocal{E}\otimes\Lambda^{2}T^{*}M$.
A Hermitian–Yang–Mills connection is a unitary connection $\nabla$ for the Hermitian metric $h$ which satisfies $F_{\nabla}^{0,2}=0$ i.e. vanishing of the curvature and the trace $tr$ of the curvature satisfies the following relation $tr(F_{\nabla})=\kappa Id{E}.$
Given a Kähler metric on $M$, define the operation $tr:\Gamma(End\pazocal{E}\otimes T_{1,1}^{*}M)\to\Gamma(End\pazocal{E})$, where $tr(F)=\sum g^{j\bar{k}}F_{aj\bar{k}}^{b}$.
We have a Hermitian–Yang–Mills if there exists a Hermitian metric $h$ for which the Hermitian curvature $F$ satisfies $tr(F)=\kappa Id$. If we take $\pazocal{E}=TM$ to be the holomorphic tangent bundle and $h=g$ then the condition above becomes the familiar Kähler-Einstein condition (definition from Uhlenbeck–Yau).